OpenStudy (icalibear):

Find the rate of change for the equation 5x + 3y = -2? PLEASE HELP :'(

OpenStudy (anonymous):

to find the rate of change of this , just differentiate with respect to x

OpenStudy (anonymous):

d(5x)/dx + d(3y)/dx = d(-2)/dx 5dx/dx+3dy/dx=0 5(1)+3dy/dx=0 5+3dy/dx=0 subtract 5 both sides 3dy/dx=-5 divide 3 both sides dy/dx=-5/3

OpenStudy (anonymous):

the rate of change is the slope of this line.